OpenStudy (anonymous):

What is the height of a triangle that has an area of 60 yd2 and a base with a length of 12 yd? A. 0.1 yd B. 2.5 yd C. 5 yd D. 10 yd

OpenStudy (cwrw238):

area = 1/2 * height * base 60 = 1/2 * h * 12 solve for h

OpenStudy (anonymous):

i dont understand

OpenStudy (cwrw238):

the formula for the area of a triangle is in my first line i substituted area = 60 and base = 12 into this equation h = height can u solve it?

OpenStudy (cwrw238):

ok lets simplify the equation right side = 1/2 * 12 * h what is 1/2 * 12?

OpenStudy (cwrw238):

what is a half of 12?

OpenStudy (anonymous):

6

OpenStudy (cwrw238):

jenny - i'm sure you know what 12 divided by 2 is....